OpenStudy (anonymous):

Find two consecutive odd integers whose sum is 72. Which equation would be used to solve this problem? (HINT: The equation has been simplified some.) 2x = 72 2x + 2 = 72 x + 2 = 72 2x + 1 = 72

OpenStudy (anonymous):

the first one ?

OpenStudy (shane_b):

Since x must be odd then: x + x + 2=72 so 2x+2=72 would be correct

OpenStudy (phi):

the odd integers go 1,3,5,7,.... pick one (anyone) say 3 the next one is 5, or 2 more so I would say call the first odd integer n and the next one is 2 up, or n+2 now add them together: n+(n+2) = 2n+2 they add up to 72 2n+2= 72

OpenStudy (anonymous):

second one ?

OpenStudy (anonymous):

yepp second one :)

OpenStudy (phi):

yes, x or n does not matter. It just means unknown

OpenStudy (anonymous):

that was my second choice

OpenStudy (phi):

If you make it your first choice you will get a better score on the test.

OpenStudy (phi):

btw, notice if you solve the first equation, you get x=36 and 36 is not odd, so that is a clue it's the wrong equation.